rssLink RSS for all categories
 
icon_red
icon_green
icon_red
icon_red
icon_blue
icon_green
icon_green
icon_red
icon_red
icon_red
icon_orange
icon_green
icon_green
icon_green
icon_green
icon_blue
icon_red
icon_orange
icon_red
icon_red
icon_red
icon_red
icon_green
icon_red
icon_red
icon_red
icon_red
icon_orange
icon_green
 

FS#378 — FS#4453 — Optic HS th1-1-

Attached to Project— Network
Incident
Whole Network
CLOSED
100%
Regarding the links' migration (Roubaix's side) task #4440, one of the optics has been disconnected at Paris's side:

Optical Optical
Temperature Voltage Current Tx Power Rx Power
Port (Celsius) (Volts) (mA) (dBm) (dBm)
------- ----------- ------- -------- -------- --------
Te5/2 31.7 0.00 1.1 -- -13.6 -- -2.2

More power output. A technician intervened to replace by a spare.

Date:  Wednesday, 11 August 2010, 02:28AM
Reason for closing:  Done
Additional comments about closing:  The optic has been replaced around 12:30.